i need to solve this problem by a system of 3 equations: A piggy bank contains 40 coins consisting of nickels, dimes, and quarters. the sum of the number of dimes and quarters is 10 coins less than the number of nickels. if the number of quarters and nickels were reversed, with the same amount of dimes, the value of the coins would be $7.50. find the number of each type of coin that is in the piggy bank.
i tried these three equations with no luck:
n+d+q=40
d+q=n-10
10d+25q+5n=750

Answer by stanbon(75887)

i need to solve this problem by a system of 3 equations: A piggy bank contains 40 coins consisting of nickels, dimes, and quarters. the sum of the number of dimes and quarters is 10 coins less than the number of nickels. if the number of quarters and nickels were reversed, with the same amount of dimes, the value of the coins would be $7.50. find the number of each type of coin that is in the piggy bank.
i tried these three equations with no luck:
n+d+q=40
d+q=n-10
10d+25q+5n=750
--------------------------
Change your 3rd equation to
25n + 10d + 5q = 750
----
Gather the equations:
n + d + q = 40
n - d - q = 10
25n+10d +5q = 750
-------------------------
I used a matrix method and got:
n = 25 (# of nickels)
d = 10 (# of dimes)
q = 5 (# of quarters)
============================
Cheers,
Stan H.